### v2

# Blow-up for a Nonlocal Diffusion Equation with Time Regularly Varying Nonlinearity and Forcing

Abstract.
We investigate the Cauchy problem for a semilinear parabolic equation driven by a mixed local–nonlocal diffusion operator of the form
$\partial_{t}u-(\Delta-(-\Delta)^{\mathsf{s}})u=\mathsf{h}(t)|x|^{-b}|u|^{p}+t^{\varrho}\mathbf{w}(x),\qquad(x,t)\in\mathbb{R}^{N}\times(0,\infty),$
where $\mathsf{s}\in(0,1)$, $p>1$, $b\geq 0$, and $\varrho>-1$. The function $\mathsf{h}(t)$ is assumed to belong to the generalized class of regularly varying functions, while $\mathbf{w}$ is a prescribed spatial source. We first revisit the unforced case and establish sharp blow-up and global existence criteria in terms of the critical Fujita exponent, thereby extending earlier results to the wider class of time-dependent coefficients. For the forced problem, we derive nonexistence of global weak solutions under suitable growth conditions on $\mathsf{h}$ and integrability assumptions on $\mathbf{w}$. Furthermore, we provide sufficient smallness conditions on the initial data and the forcing term ensuring global-in-time mild solutions. Our analysis combines semigroup estimates for the mixed operator, test function methods, and asymptotic properties of regularly varying functions. To our knowledge, this is the first study addressing blow-up phenomena for nonlinear diffusion equations with such a class of time-dependent coefficients.
Key words and phrases:
Non-local parabolic equation, finite time blow-up global existence, critical exponent, forcing term, regularly varying functions.
2020 Mathematics Subject Classification:
35A01, 35K57, 35K58, 35C15, 35B44, 35R11, 26A12.
1. Introduction and main results
In this paper, we study the blow-up phenomenon for solutions to the following mixed local–nonlocal diffusion equation:
$\begin{cases}\partial_{t}u(x,t)-\mathscr{L}u(x,t)=\mathsf{h}(t)\,|x|^{-b}|u(x,t)|^{p}+t^{\varrho}\,\mathbf{w}(x),\\ u(x,0)=u_{0}(x),\end{cases}$ (1.1)
where $(x,t)\in\mathbb{R}^{N}\times(0,\infty)$, the parameters satisfy $p>1$, $b\geq 0$, and $\varrho>-1$. The time-dependent coefficient $\mathsf{h}\colon(0,\infty)\to(0,\infty)$ is a given continuous function and $\mathbf{w}\colon\mathbb{R}^{N}\to\mathbb{R}$ is a prescribed spatial function. The diffusion operator $\mathscr{L}$ is defined by
$\mathscr{L}:=\Delta-(-\Delta)^{\mathsf{s}},\quad\text{with }\mathsf{s}\in(0,1),$
and models a combination of classical local diffusion (via the Laplace operator $\Delta$) and nonlocal diffusion (via the fractional Laplace operator $(-\Delta)^{\mathsf{s}}$).
The mixed local-nonlocal operator $\mathscr{L}$ combines the classical Laplacian $\Delta$, a local second-order differential operator, with the fractional Laplacian $(-\Delta)^{s}$, which is nonlocal. The classical Laplacian models standard diffusion processes such as Brownian motion, while the fractional Laplacian accounts for anomalous diffusion characterized by long-range jumps, as in Lévy flights [21, 49]. The operator $\mathscr{L}$ thus describes a competition between local and non-local diffusion, making it suitable for modeling phenomena where both short- and long-range interactions coexist.
It is worth noticing that the fractional Laplacian arises naturally in the theory of stochastic processes, particularly in connection with symmetric $\alpha$-stable Lévy processes. These processes, which generalize Brownian motion by allowing for jumps, are characterized by independent and stationary increments, and their paths exhibit discontinuities.
Furthermore, symmetric $\alpha$-stable Lévy processes can be constructed by subordinating a Brownian motion with an increasing Lévy process, known as a subordinator. This probabilistic perspective leads to natural connections between nonlocal evolution equations and stochastic processes. For further details and foundational results, we refer to the works of Applebaum [3], Bertoin [8], and Bogdan et al. [12], among others.
Nonlocal models have gained significant attention as robust alternatives to classical partial differential equations (PDEs), especially when local formulations fail to accurately describe phenomena involving multiscale interactions or anomalous transport. A wide range of physical and engineering systems exhibit intrinsic nonlocality and hierarchical structures that render classical PDE-based models inadequate. Such features are prevalent in various applications, including continuum mechanics [27, 34, 48], phase transitions [6, 13, 17], corrosion processes [42], turbulent flows [4, 40, 43], and geophysical modeling [7, 37, 44, 45, 50]. These settings often require mathematical frameworks that incorporate long-range interactions or fractional-order operators to capture the underlying dynamics more faithfully.
From a mathematical point of view, the operator $\mathscr{L}$ is of significant interest due to the interplay between local non-local dynamics. It presents new challenges in analysis, including the study of regularity, spectral properties, and maximum principles. Equations involving $\mathscr{L}$, such as
$\partial_{t}u=\mathscr{L}u+f(u),$ (1.2)
serve as a framework for reaction-diffusion models with mixed diffusion.
The investigation of blow-up phenomena for equation (1.2) dates back to  [47], which considered the case of the pure fractional Laplacian $\mathscr{L}=-(-\Delta)^{\mathsf{s}}$ along with general nonlinearities. Since then, various aspects of the purely fractional setting, featuring different types of nonlinearities and alternative analytical techniques, have been further developed in [22, 25, 24].
In recent work, Biagi, Punzo, and Vecchi [9], and subsequently Del Pezzo and Ferreira [20], investigated Fujita-type phenomena for equation (1.2) involving a power-type nonlinearity $f(u)=u^{p}$. They identified the critical Fujita exponent as $1+\frac{2\mathsf{s}}{N}$, which marks the threshold between global existence and finite-time blow-up. Their analysis, which relies on the classical Kaplan eigenfunction method [29], reveals an intriguing result: the critical exponent $1+\frac{2\mathsf{s}}{N}$ coincides exactly with that of the purely fractional Laplacian case. This indicates that the presence of a local diffusion term does not alter the fundamental blow-up behavior governed by the fractional component. In essence, the mixed local-nonlocal operator preserves the same criticality as the fractional Laplacian in determining the long-time dynamics of solutions.
Regarding the existence of global solutions, the strategy in [9] relies on an approximation scheme to build suitable solutions step by step. On the other hand, the approach in [20] is based on the explicit construction of a global supersolution, which serves as an upper barrier to control the behavior of solutions over time.
Recently, the problem (1.1) in the case $b=0$ and without a forcing term was studied in [15], where the authors considered more general nonlinearities beyond the standard power-type case. In particular, they explored the initial value problem
$\begin{cases}\partial_{t}u-\mathscr{L}u=\mathsf{h}(t)\,u^{p}&\text{in }\mathbb{R}^{N}\times(0,\infty),\\ u(x,0)=u_{0}(x)\geq 0&\text{in }\mathbb{R}^{N},\end{cases}$ (1.3)
where $\mathsf{h}\in C([0,\infty))$ is a nonnegative function. The main result presented in [15, Theorem 6], when adapted to the framework of equation (1.3), can be summarized as follows:
[…]
    Suppose $v_{0}\in L^{1}\cap L^{\infty}$ is nonnegative. If the following integral condition holds:
    $\int\limits_{0}^{\infty}\mathsf{h}(\tau)\|e^{\tau\mathscr{L}}v_{0}\|_{\infty}^{p-1}\,d\tau<1,$ (1.4)
    then there exists a constant $\delta>0$ such that the solution to (1.3) with initial data $u_{0}=\delta\,v_{0}$ exists globally in time.
[…]
    On the other hand, if $u_{0}\in L^{1}\cap L^{\infty}$ is nontrivial and nonnegative, and there exists some $t_{0}>0$ such that
    $(p-1)\,\|e^{t_{0}\mathscr{L}}u_{0}\|_{\infty}^{p-1}\int\limits_{0}^{t_{0}}\mathsf{h}(\tau)\,d\tau\geq 1,$ (1.5)
    then the corresponding mild solution to (1.3) blows up in finite time.
As an application of [15, Theorem 6], the authors determine the Fujita exponent for equation (1.3) under the assumption that the function $\mathsf{h}$ satisfies the asymptotic growth condition
$C_{1}t^{\gamma}\leq\mathsf{h}(t)\leq C_{2}t^{\gamma},\quad\text{for }t\gg 1,$ (1.6)
for some $\gamma>0$ and constants $C_{1},C_{2}>0$. Under this assumption, they show that the critical Fujita exponent is given by
$p_{F}=1+\frac{2\mathsf{s}(\gamma+1)}{N}.$ (1.7)
Interestingly, the growth condition (1.6) can be interpreted as saying that $\mathsf{h}\in\mathcal{M}(\gamma)$, where the class $\mathcal{M}(\gamma)$ is given in the Definition A.2 below. As explained in Appendix A, this class serves as a natural extension of the classical class of regularly varying functions.
Before presenting our main results concerning equation (1.1), we first clarify the notions of weak and mild solutions.
Definition 1.1.
We say that a function $u(x,t)$ is a global weak solution of (1.1) if it satisfies the following conditions:
$u_{0}\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{N}),\quad\mathsf{h}(t)\,|x|^{-b}|u|^{p}\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{N}\times(0,\infty)),$
and for every test function $\psi\in C^{\infty}_{0}(\mathbb{R}^{N}\times(0,\infty))$, the identity
$\begin{split}\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}u(-\partial_{t}\psi-\mathscr{L}\psi)\,dx\,dt&=\int\limits_{\mathbb{R}^{N}}u_{0}(x)\,\psi(x,0)\,dx+\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}\mathsf{h}(t)\,|x|^{-b}|u|^{p}\psi\,dx\,dt\\ &\quad+\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}t^{\varrho}\,\mathbf{w}(x)\,\psi\,dx\,dt\end{split}$ (1.8)
holds.
Alternatively, equation (1.1) can be expressed in its Duhamel form as
$u(x,t)\;=\;e^{t\mathscr{L}}u_{0}+\int_{0}^{t}\mathsf{h}(s)\,e^{(t-s)\mathscr{L}}\!\left(|\cdot|^{-b}\,|u(s)|^{p}\right)ds+\int_{0}^{t}s^{\varrho}\,e^{(t-s)\mathscr{L}}\mathbf{w}(\cdot)\,ds,$ (1.9)
where $e^{t\mathscr{L}}$ denotes the semigroup generated by the mixed local–nonlocal operator $\mathscr{L}$ (see Section 2 for details). A function $u$ that satisfies (1.9) is referred to as a mild solution of (1.1).
Our main result on equation (1.3) extends the scope of [15, Corollary 7] by allowing the function $\mathsf{h}$ to belong to the broader class $\mathcal{M}(\gamma)$, for any $\gamma>-1$.
Theorem 1.1.
Assume that $\mathsf{h}\in\mathcal{M}(\gamma)$ for some $\gamma>-1$. Then the following holds:
  1. (i)
    If $p<1+\dfrac{2\mathsf{s}(\gamma+1)}{N}$, then every nonnegative solution of (1.3) blows up in finite time.
  2. (ii)
    If $p>1+\dfrac{2\mathsf{s}(\gamma+1)}{N}$, then equation (1.3) admits a global-in-time solution for sufficiently small initial data.
Remark 1.1.
  1. (i)
    As will become clear later, the proof of Theorem 1.1 relies on [15, Theorem 6] together with key properties of the function class $\mathcal{M}(\gamma)$, which are discussed in Appendix A.
  2. (ii)
    Examples of functions belonging to $\mathcal{M}(\gamma)$ include
    $t^{\gamma}\log(1+t),\quad t^{\gamma}(2+\sin(\log t)),\quad t^{\gamma}\exp\left(\sqrt{|\log t|}\right).$
    More generally, one can consider functions of the form $\mathsf{h}(t)=t^{\gamma}\ell(t)$, where $\ell$ is slowly varying at infinity in the sense of  (A.2).
Our next result addresses the nonexistence of global solutions to (1.1) in the presence of a forcing term. The precise statement is as follows:
Theorem 1.2.
Suppose that the function $\mathsf{h}$ is given by
$\mathsf{h}(t)=t^{\gamma}\,\ell(t),$ (1.10)
where $\gamma>-1$ and $\ell:(0,\infty)\to(0,\infty)$ is slowly varying at infinity. Assume further that ${\mathbf{w}}\in C_{0}(\mathbb{R}^{N})\cap L^{1}(\mathbb{R}^{N})$ satisfies
$\int\limits_{\mathbb{R}^{N}}{\mathbf{w}}(x)\,dx>0.$ (1.11)
  1. (i)
    If $\varrho\leq 0$, $0\leq\dfrac{b}{1+\gamma}<2\mathsf{s}<N$, and
    $1<p<p^{*}\vcentcolon=\frac{N-b-2\mathsf{s}(\varrho-\gamma)}{\,N-2\mathsf{s}(\varrho+1)},$ (1.12)
    then problem (1.1) admits no global weak solution in the sense of Definition 1.1.
  2. (ii)
    If $\varrho>0$ and $b,\gamma\geq 0$, then the same conclusion holds for every $p>1$.
Remark 1.2.
  1. (i)
    The case $\mathsf{h}=1$ with $b=\varrho=0$ was recently studied in [33].
  2. (ii)
    Condition $\dfrac{b}{1+\gamma}<2\mathsf{s}$ guarantees that the exponent $p^{*}$ in (1.12) satisfies $p^{*}>1$.
  3. (iii)
    For related results in the case $\mathsf{s}=1$, we refer the reader to [1, 5, 28, 35]. In particular, when $b=\gamma=0$ and $\varrho\in(-1,0)$, one has
    $p^{*}=\frac{N-2\varrho}{\,N-2\varrho-2},$
    in agreement with [28, Theorem 1.1, (1.8)].
  4. (iv)
    The particular case where $\mathsf{h}(t)=1$ and $\mathsf{s}\geq 1$ is an integer was previously studied in [36].
  5. (v)
    One of the main novelties of this work, beyond the use of a mixed local-nonlocal operator, is the general form (1.10) assumed for the function $\mathsf{h}$. To the best of our knowledge, this is the first time such a class of time-dependent coefficients is considered in the study of blow-up phenomena for nonlinear diffusion equations.
  6. (vi)
    As the proof will show, assumption (1.10) can be relaxed to $\mathsf{h}\in\mathcal{M}(\gamma)$
Switching now to the analysis of the global theory, we establish the following global existence result.
Theorem 1.3.
Assume that the function $\mathsf{h}$ is given by (1.10). Let $\mathsf{s}\in(0,1)$, $0\leq\dfrac{b}{1+\gamma}<2\mathsf{s}<N$, and $1<\varrho<0$. Suppose that the exponent $p$ satisfies
$p\;\geq\;p^{*},$ (1.13)
where $p^{*}$ is defined in (1.12). Define the critical exponents
$p_{c}$ $:=\frac{N(p-1)}{2\mathsf{s}(1+\gamma)-b},$ (1.14)
$q_{c}$ $:=\frac{Np_{c}}{N+2\mathsf{s}(\varrho+1)p_{c}}.$ (1.15)
Then, there exists a constant $\epsilon>0$ such that, if the initial data $u_{0}$ and the external force term $\mathbf{w}$ satisfy
$\|u_{0}\|_{L^{p_{c}}(\mathbb{R}^{N})}+\|\mathbf{w}\|_{L^{q_{c}}(\mathbb{R}^{N})}\;<\;\epsilon,$ (1.16)
the problem (1.1) admits a global-in-time mild solution $u$.
Remark 1.3.
  1. (i)
    Observe that, using (1.14) and (1.15) together with the condition $\varrho>-1$, we obtain
    $\frac{1}{q_{c}}=\frac{1}{p_{c}}+\frac{2\mathsf{s}(\varrho+1)}{N}>\frac{1}{p_{c}}.$
  2. (ii)
    Results similar to Theorem 1.3 have been obtained in [33, 28, 35, 36] and the references therein.
The structure of the article is as follows. In Section 2, we introduce the notation used throughout the paper and present several auxiliary results and estimates. Section 3 is devoted to the study of the unforced problem (1.3), where we provide the proof of Theorem 1.1. In Section 4, we address the main problem (1.1) and establish Theorems 1.2 and 1.3. Concluding remarks and directions for future research are given in Section 5. Finally, Appendix A contains a brief overview of regularly varying functions together with several useful estimates employed in our analysis.
Throughout the remainder of the article, the constant $C>0$ may vary from line to line. We write $X\lesssim Y$ or $Y\gtrsim X$ to indicate the inequality $X\leq CY$ for some constant $C>0$. The Lebesgue norm $\|\cdot\|_{L^{r}(\mathbb{R}^{N})}$ is denoted by $\|\cdot\|_{r}$ for $1\leq r\leq\infty$.
2. Useful tools & Auxiliary results
In this section, we introduce the notation used throughout the paper and present several auxiliary results and estimates.
The fractional Laplacian operator $(-\Delta)^{\mathsf{s}}$ with $\mathsf{s}\in(0,1)$ generates a semigroup $\{e^{-t(-\Delta)^{\mathsf{s}}}\}_{t\geq 0}$, whose kernel $\mathscr{E}_{\mathsf{s}}$ is smooth, radial, and satisfies the scaling property
$\mathscr{E}_{\mathsf{s}}(x,t)=t^{-\frac{N}{2\mathsf{s}}}\,\mathscr{K}_{\mathsf{s}}\!\left(t^{-\frac{1}{2\mathsf{s}}}x\right),$ (2.1)
where the profile function $\mathscr{K}_{\mathsf{s}}$ is given by the Fourier integral
$\mathscr{K}_{\mathsf{s}}(x)=(2\pi)^{-N/2}\int\limits_{\mathbb{R}^{N}}e^{ix\cdot\xi}e^{-|\xi|^{2\mathsf{s}}}\,d\xi.$ (2.2)
Explicit formulas for $\mathscr{E}_{\mathsf{s}}$ are available in two important cases:
  • For $\mathsf{s}=1$ (standard heat kernel):
    $\mathscr{E}_{1}(x,t)=(4\pi t)^{-N/2}e^{-\tfrac{|x|^{2}}{4t}},\qquad\mathscr{K}_{1}(x)=(4\pi)^{-N/2}e^{-\tfrac{|x|^{2}}{4}}.$ (2.3)
  • For $\mathsf{s}=\tfrac{1}{2}$ (Poisson kernel):
    $\mathscr{E}_{1/2}(x,t)=\frac{\Gamma\!\left(\tfrac{N+1}{2}\right)t}{\pi^{\frac{N+1}{2}}(t^{2}+|x|^{2})^{\frac{N+1}{2}}},\qquad\mathscr{K}_{1/2}(x)=\frac{\Gamma\!\left(\tfrac{N+1}{2}\right)}{\pi^{\frac{N+1}{2}}(1+|x|^{2})^{\frac{N+1}{2}}}.$ (2.4)
For general $\mathsf{s}\in(0,1)$, while no explicit representation is known, the following positivity estimate holds.
Lemma 2.1.
Let $N\geq 1$ and $\mathsf{s}\in(0,1)$. Then the profile function $\mathscr{K}_{\mathsf{s}}$ satisfies
$(1+|x|)^{-N-2\mathsf{s}}\,\lesssim\,\mathscr{K}_{\mathsf{s}}(x)\,\lesssim\,(1+|x|)^{-N-2\mathsf{s}},\qquad x\in\mathbb{R}^{N}.$ (2.5)
In particular, $\mathscr{K}_{\mathsf{s}}\in L^{p}(\mathbb{R}^{N})$ for all $1\leq p\leq\infty$.
The proof appears in [2, p. 395], while the positivity result was first stated without proof in [11, p. 263]. A detailed argument can also be found in [11, Theorem 2.1].
The operator $\mathscr{L}=\Delta-(-\Delta)^{\mathsf{s}}$ generates a strongly continuous contraction semigroup $\{e^{t\mathscr{L}}\}_{t\geq 0}$ on $L^{2}(\mathbb{R}^{N})$, where each operator $e^{t\mathscr{L}}$ is given by convolution with the fundamental solution $\mathbf{E}_{\mathsf{s}}(t)$. This fundamental solution $\mathbf{E}_{\mathsf{s}}(x,t)$ solves the evolution equation
$\partial_{t}u(x,t)=\mathscr{L}u(x,t),\qquad(x,t)\in\mathbb{R}^{N}\times(0,\infty),$ (2.6)
with Dirac mass as initial data. It can be expressed as the convolution of the classical heat kernel $\mathscr{E}_{1}(x,t)=(4\pi t)^{-N/2}e^{-\frac{|x|^{2}}{4t}}$ and the fractional heat kernel $\mathscr{E}_{\mathsf{s}}(x,t)$ from (2.1).
The fundamental solution $\mathbf{E}_{\mathsf{s}}(x,t)$ enjoys several important properties (see, e.g., [32] and [9]).
Lemma 2.2.
Let $\mathsf{s}\in(0,1)$. Then the following hold:
  1. (i)
    Regularity and positivity: $\mathbf{E}_{\mathsf{s}}\in C^{\infty}(\mathbb{R}^{N}\times(0,\infty))$ and $\mathbf{E}_{\mathsf{s}}(x,t)\geq 0$ for all $(x,t)\in\mathbb{R}^{N}\times(0,\infty)$.
  2. (ii)
    Mass conservation: The kernel preserves total mass:
    $\int\limits_{\mathbb{R}^{N}}\mathbf{E}_{\mathsf{s}}(x,t)\,dx=1,\qquad t>0.$ (2.7)
  3. (iii)
    Smoothing estimates: For any $\varphi\in L^{r}(\mathbb{R}^{N})$ and $1\leq r\leq q\leq\infty$, we have
    $\|\mathbf{E}_{\mathsf{s}}(t)\ast\varphi\|_{q}\;\leq\;C\min\!\left\{t^{-\frac{N}{2}\left(\frac{1}{r}-\frac{1}{q}\right)},\;t^{-\frac{N}{2\mathsf{s}}\left(\frac{1}{r}-\frac{1}{q}\right)}\right\}\|\varphi\|_{r},\qquad t>0.$ (2.8)
As a consequence of Lemma 2.2, we obtain the following bounds for the semigroup $e^{t\mathscr{L}}$.
Lemma 2.3.
Let $\varphi\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ and $t>0$. Then:
  1. (i)
    Contractivity and decay:
    $\|e^{t\mathscr{L}}\varphi\|_{1}\leq\|\varphi\|_{1},\qquad\|e^{t\mathscr{L}}\varphi\|_{\infty}\leq\|\varphi\|_{\infty},\qquad\|e^{t\mathscr{L}}\varphi\|_{\infty}\lesssim t^{-\frac{N}{2\mathsf{s}}}.$
  2. (ii)
    Lower bound: If, in addition, $\varphi\geq 0$ and $\varphi\not\equiv 0$, then
    $\|e^{t\mathscr{L}}\varphi\|_{\infty}\;\gtrsim\;t^{-\frac{N}{2\mathsf{s}}}.$ (2.9)
Remark 2.1.
A proof of Lemma 2.3 can be found, for instance, in [15].
The next lemma, adapted from [41], provides a useful convexity inequality for the fractional Laplacian.
Lemma 2.4.
Let $\mathsf{s}\in(0,1)$, let $G\in C^{2}(\mathbb{R},\mathbb{R})$ be convex, and let $\phi:\mathbb{R}^{N}\to\mathbb{R}$ be smooth and compactly supported. Then the following inequality holds:
$(-\Delta)^{\mathsf{s}}\bigl{[}G(\phi)\bigr{]}\;\leq\;G^{\prime}(\phi)\,(-\Delta)^{\mathsf{s}}\phi,\qquad\text{in }\mathbb{R}^{N}.$ (2.10)
3. The unforced problem
In this section, we provide the proof of Theorem 1.1, which addresses the unforced problem (1.3).
  1. (i)
    Suppose that $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}.$ According to [15, Theorem 6], it suffices to show that (1.5) holds for some $t_{0}>0$. By Lemma 2.3, we have
    $(p-1)\,\|e^{t\mathscr{L}}u_{0}\|_{\infty}^{p-1}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\geq\;C\,t^{-\frac{N(p-1)}{2\mathsf{s}}}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau.$ (3.1)
    Moreover, by (A.14), there exists a slowly varying function $\ell$ such that, for all sufficiently large $t>0$,
    $\mathsf{h}(t)\gtrsim t^{\gamma}\ell(t).$ (3.2)
    Consequently, for $t>0$ large enough,
    $\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\geq\;\int\limits_{t/2}^{t}\mathsf{h}(\tau)\,d\tau\;\gtrsim\;t^{\gamma+1}.$ (3.3)
    Inserting (3.3) into (3.1) yields
    $(p-1)\,\|e^{t\mathscr{L}}u_{0}\|_{\infty}^{p-1}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\gtrsim\;t^{\gamma+1-\frac{N(p-1)}{2\mathsf{s}}}.$ (3.4)
    Since $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}$, we have $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}>0.$ Therefore, one can choose $t_{0}>0$ sufficiently large so that (1.5) holds. This completes the proof of the first part of Theorem 1.1.
  2. (ii)
    Assume now that $p>1+\frac{2\mathsf{s}(\gamma+1)}{N}.$ We will prove that condition (1.4) holds, which guarantees the global existence of solutions. For $t_{0}>0$ (to be chosen sufficiently large later), applying Lemma 2.3 together with (A.14), we obtain
    $\begin{split}\int\limits_{0}^{\infty}\mathsf{h}(\tau)\,\|e^{\tau\mathscr{L}}v_{0}\|_{\infty}^{\,p-1}\,d\tau&\leq\left(\int\limits_{0}^{t_{0}}\mathsf{h}(\tau)\,d\tau\right)\|v_{0}\|_{\infty}^{\,p-1}+C\int\limits_{t_{0}}^{\infty}\mathsf{h}(\tau)\,\tau^{-\frac{N(p-1)}{2\mathsf{s}}}\,d\tau\\ &\leq\left(\int\limits_{0}^{t_{0}}\mathsf{h}(\tau)\,d\tau\right)\|v_{0}\|_{\infty}^{\,p-1}+C\int\limits_{t_{0}}^{\infty}\ell_{1}(\tau)\,\tau^{\gamma-\frac{N(p-1)}{2\mathsf{s}}}\,d\tau,\end{split}$ (3.5)
    where $\ell_{1}$ is a slowly varying function provided by Theorem A.4. Since $\gamma-\frac{N(p-1)}{2\mathsf{s}}<-1,$ Theorem A.3 yields
    $\int\limits_{0}^{\infty}\mathsf{h}(\tau)\,\|e^{\tau\mathscr{L}}v_{0}\|_{\infty}^{\,p-1}\,d\tau\leq\left(\int\limits_{0}^{t_{0}}\mathsf{h}(\tau)\,d\tau\right)\|v_{0}\|_{\infty}^{\,p-1}+C\,t_{0}^{\,\gamma+1-\frac{N(p-1)}{2\mathsf{s}}}.$ (3.6)
    Finally, since $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}<0$, the desired result follows from [15, Theorem 6], by first choosing $t_{0}>0$ sufficiently large and then taking $\|v_{0}\|_{\infty}$ small enough.
4. The forced problem
4.1. Proof of Theorem 1.2
Let $\mathbf{w}\in C_{0}(\mathbb{R}^{N})\cap L^{1}(\mathbb{R}^{N})$ be such that $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx>0$. Suppose further that (1.10) holds with $\gamma>-1$. We proceed by contradiction, assuming that Problem (1.1) possesses a global weak solution in the sense of Definition 1.1.
  1. (i)
    Here we assume that
    $\varrho\leq 0,\quad 0\leq\frac{b}{1+\gamma}<2\mathsf{s}<N,$
    and that condition (1.12) is satisfied. In this setting, we shall employ a test function method, which is commonly used in this context (see, e.g., [33, 26, 25, 24, 28, 35, 36, 38]).
    Let $\eta,\phi\in C^{\infty}_{0}([0,\infty))$ be cut-off functions such that $0\leq\eta,\phi\leq 1$ and
    $\eta(r)=\begin{cases}1,&\text{if }\tfrac{1}{2}\leq r\leq\tfrac{3}{4},\\[4.0pt] 0,&\text{if }r\in[0,\tfrac{1}{4}]\cup[\tfrac{4}{5},\infty),\end{cases}\qquad\phi(r)=\begin{cases}1,&\text{if }0\leq r\leq 1,\\[4.0pt] 0,&\text{if }r\geq 2.\end{cases}$
    For sufficiently large $R>0$, we define the test function
    $\psi_{R}(x,t)=\phi^{\mathbf{m}}\!\left(\tfrac{|x|}{R}\right)\,\eta^{\mathbf{m}}\!\left(\tfrac{t}{R^{2\mathsf{s}}}\right),\qquad\text{where }\mathbf{m}=\frac{2p}{p-1}>2.$ (4.1)
    Since $u$ is a global weak solution of (1.1) and
    $\int\limits_{\mathbb{R}^{N}}\psi_{R}(x,0)\,u_{0}(x)\,dx=0,$
    the weak formulation (1.8) implies that
    $\begin{split}\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}h(t)|x|^{-b}|u|^{p}\psi_{R}\,dx\,dt&+\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}t^{\varrho}\mathbf{w}(x)\psi_{R}\,dx\,dt\\ &\leq\underbrace{\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}|u||\partial_{t}\psi_{R}|\,dx\,dt}_{\mathcal{I}}+\underbrace{\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}|u||\mathscr{L}\psi_{R}|\,dx\,dt}_{\mathcal{J}}.\end{split}$ (4.2)
    Applying the $\varepsilon$-Young inequality, we obtain the estimates
    $\mathcal{I}$ $\leq\frac{1}{4}\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}h(t)|x|^{-b}|u|^{p}\psi_{R}\,dx\,dt+C\underbrace{\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}h(t)^{-\frac{1}{p-1}}|x|^{\frac{b}{p-1}}\psi_{R}^{-\frac{1}{p-1}}|\partial_{t}\psi_{R}|^{\frac{p}{p-1}}\,dx\,dt}_{\mathcal{I}_{1}},$ (4.3)
    $\mathcal{J}$ $\leq\frac{1}{4}\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}\mathsf{h}(t)|x|^{-b}|u|^{p}\psi_{R}\,dx\,dt+C\underbrace{\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}\mathsf{h}(t)^{-\frac{1}{p-1}}|x|^{\frac{b}{p-1}}\psi_{R}^{-\frac{1}{p-1}}|\mathscr{L}\psi_{R}|^{\frac{p}{p-1}}\,dx\,dt}_{\mathcal{J}_{1}}.$ (4.4)
    Exploiting the support properties of the cut-off functions $\eta$ and $\phi$, we estimate the term $\mathcal{I}_{1}$ as
    $\begin{split}\mathcal{I}_{1}&\lesssim R^{-\frac{2p\mathsf{s}}{p-1}}\left(\int\limits_{\frac{R^{2\mathsf{s}}}{4}}^{\frac{4R^{2\mathsf{s}}}{5}}\mathsf{h}(t)^{-\frac{1}{p-1}}|\eta^{\prime}|^{\frac{p}{p-1}}|\eta|^{\mathbf{m}-\frac{p}{p-1}}\,dt\right)\left(\int\limits_{\{|x|\leq 2R\}}|x|^{\frac{b}{p-1}}\phi^{\mathbf{m}}(x)\,dx\right)\\ &\lesssim R^{N+\frac{b-2\mathsf{s}}{p-1}}\left(\int\limits_{\frac{1}{4}}^{\frac{4}{5}}\Big{(}\mathsf{h}(R^{2\mathsf{s}}\tau)\Big{)}^{-\frac{1}{p-1}}\,d\tau\right).\end{split}$ (4.5)
    To estimate the second term $\mathcal{J}_{1}$, we again use the support properties of $\phi$, together with $0\leq\phi\leq 1$ and (2.10), to first obtain, for $R\geq 1$,
    $|\Delta\phi^{\mathbf{m}}|\lesssim R^{-2}\lesssim R^{-2\mathsf{s}},\qquad|(-\Delta)^{\mathsf{s}}\phi^{\mathbf{m}}|\lesssim R^{-2\mathsf{s}}.$ (4.6)
    This implies
    $|\mathscr{L}\phi^{\mathbf{m}}|\lesssim R^{-2\mathsf{s}},\qquad R\geq 1.$ (4.7)
    Using (4.7) and arguing as in the case of $\mathcal{I}_{1}$, we arrive at
    $\mathcal{J}_{1}\lesssim R^{N+\frac{b-2\mathsf{s}}{p-1}}\left(\int\limits_{\frac{1}{4}}^{\frac{4}{5}}\Big{(}\mathsf{h}(R^{2\mathsf{s}}\tau)\Big{)}^{-\frac{1}{p-1}}\,d\tau\right),\qquad R\geq 1.$ (4.8)
    Combining (4.2) with (4.3), (4.4), (4.5), and (4.8), we obtain
    $\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}t^{\varrho}\mathbf{w}(x)\psi_{R}(x,t)\,dx\,dt\lesssim R^{N+\frac{b-2\mathsf{s}}{p-1}}\left(\int\limits_{\frac{1}{4}}^{\frac{4}{5}}\Big{(}\mathsf{h}(R^{2\mathsf{s}}\tau)\Big{)}^{-\frac{1}{p-1}}\,d\tau\right),\qquad R\geq 1.$ (4.9)
    On the other hand, since $\mathbf{w}\in L^{1}$ and $\int\limits\mathbf{w}(x)\,dx>0$, Lebesgue’s theorem ensures that, for sufficiently large $R\geq 1$,
    $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,\phi^{\mathbf{m}}\!\left(\tfrac{|x|}{R}\right)dx\;\geq\;\tfrac{1}{2}\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx.$ (4.10)
    Therefore, the left-hand side of (4.9) can be bounded from below, for sufficiently large $R\geq 1$, as
    $\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{N}}t^{\varrho}\mathbf{w}(x)\psi_{R}(x,t)\,dx\,dt\;\gtrsim\;R^{2\mathsf{s}(\varrho+1)}\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx.$ (4.11)
    Plugging (4.11) into (4.9), using the expression of $\mathsf{h}(t)$ in (1.10), and invoking Proposition A.1, we infer
    $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx\;\lesssim\;R^{\,N-2\mathsf{s}(\varrho+1)+\frac{b-2\mathsf{s}(1+\gamma)}{p-1}}\Big{(}\ell(R^{2\mathsf{s}})\Big{)}^{-\frac{1}{p-1}}.$ (4.12)
    Finally, thanks to Lemma A.1 and the fact that
    $N-2\mathsf{s}(\varrho+1)+\frac{b-2\mathsf{s}(1+\gamma)}{p-1}<0,$
    we deduce, by letting $R\to\infty$ in (4.12), that
    $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx\;\leq\;0,$
    which is a contradiction. Hence, the proof of the first part of Theorem 1.2 is complete.
  2. (ii)
    Assume now that $\varrho>0$ and $b,\gamma\geq 0$. We adapt the previous approach with a slightly modified test function. More precisely, for $R,T>0$, we replace the test function defined in (4.1) by
    $\psi_{R,T}(x,t)=\phi^{\mathbf{m}}\!\left(\tfrac{|x|}{R}\right)\eta^{\mathbf{m}}\!\left(\tfrac{t}{T}\right),\qquad\text{where }\mathbf{m}=\frac{2p}{p-1}.$ (4.13)
    Proceeding as in the first part of the proof, we obtain, for $R$ sufficiently large,
    $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx\;\lesssim\;R^{\,N+\frac{b}{p-1}}\left(T^{-1-\varrho-\frac{1+\gamma}{p-1}}+T^{-\varrho-\frac{\gamma}{p-1}}R^{-\frac{2p\mathsf{s}}{p-1}}\right)\big{(}\ell(T)\big{)}^{-\frac{1}{p-1}}.$ (4.14)
    Since $\varrho>0$ and $\gamma\geq 0$, Lemma A.1 ensures that, letting $T\to\infty$ in (4.14), we obtain
    $\int\limits_{\mathbb{R}^{N}}\mathbf{w}(x)\,dx\;\leq\;0,$
    which yields a contradiction. Thus, the proof of Theorem 1.2 is completely finished.
4.2. Proof of Theorem 1.3
The proof is inspired by [16, 28, 36, 33], where a fixed-point argument is employed in a suitable complete metric space. Here, we provide only the main ingredients. Arguing as in the proof of [36, Theorem 1.7], we distinguish two cases:
$p^{*}\leq p\leq\frac{N-b+2\mathsf{s}\gamma}{N-2\mathsf{s}}\qquad\text{and}\qquad p>\frac{N-b+2\mathsf{s}\gamma}{N-2\mathsf{s}}.$
Let us consider first the case $p^{*}\leq p\leq\tfrac{N-b+2\mathsf{s}\gamma}{N-2\mathsf{s}}$. From (1.13), (1.14), (1.15), and the assumptions
$0\leq\frac{b}{1+\gamma}<2\mathsf{s}<N,\qquad-1<\varrho<0,$
we claim that there exists $r>p$ such that
$\max\left\{\,\frac{1}{p_{c}}+\frac{2\varrho\mathsf{s}}{N},\;\frac{1}{pp_{c}}\,\right\}<\frac{1}{r}<\min\left\{\,\frac{1}{p_{c}},\;\frac{N-2\mathsf{s}(\varrho+1)}{Np}\,\right\}.$ (4.15)
Such a choice is indeed possible. In fact, under the above conditions, one can readily check that all inequalities in (4.15) are satisfied. Note also that $r>p_{c}>q_{c}\geq 1$.
Next, we set
$\mu=\frac{N}{2\mathsf{s}(\gamma+1)-b}\left(\frac{1}{p_{c}}-\frac{1}{r}\right).$ (4.16)
A direct computation shows that
$0<\mu<\frac{1}{p},$ (4.17)
$\mu=\frac{N}{2\mathsf{s}(\gamma+1)-b}\left(\frac{1}{q_{c}}-\frac{1}{r}\right)-\frac{2\mathsf{s}(\varrho+1)}{2\mathsf{s}(\gamma+1)-b},$ (4.18)
and
$(1-p)\mu+1-\frac{N}{(2\mathsf{s}(\gamma+1)-b)r}(p-1)=0.$ (4.19)
We now introduce the set
$\mathbf{E}=\Bigl{\{}\,u\in L^{\infty}\bigl{(}(0,\infty);L^{r}(\mathbb{R}^{N})\bigr{)}:\;\sup_{t>0}t^{\mu}\|u(t)\|_{r}\leq\epsilon\,\Bigr{\}},$
endowed with the distance
$d(u,v)=\sup_{t>0}t^{\mu}\|u(t)-v(t)\|_{r},\qquad u,v\in\mathbf{E}.$
Then $(\mathbf{E},d)$ is a complete metric space.
Given $u\in\mathbf{E}$, define
$\Phi(u)(t)=e^{t\mathscr{L}}u_{0}+\int\limits_{0}^{t}\mathsf{h}(s)\,e^{(t-s)\mathscr{L}}\!\left(|\cdot|^{-b}|u(s)|^{p}\right)ds+\int\limits_{0}^{t}s^{\varrho}e^{(t-s)\mathscr{L}}\mathbf{w}(\cdot)\,ds.$
Using Lemma 2.3 together with (4.15)–(4.19), one easily verifies that $\Phi:\mathbf{E}\to\mathbf{E}$ is a contraction for $\epsilon$ sufficiently small. By the Picard fixed-point theorem, this yields a global solution $u\in L^{\infty}\bigl{(}(0,\infty);L^{r}(\mathbb{R}^{N})\bigr{)}$.
The reminder case $p>\tfrac{N-b+2\mathsf{s}\gamma}{N-2\mathsf{s}}$ can be handled in a similar way to [36].
5. Conclusion and Open Problems
In this paper, we analyzed the Cauchy problem for a semilinear parabolic equation involving a mixed local–nonlocal diffusion operator, a time-dependent coefficient $\mathsf{h}(t)$ taken from the generalized class of regularly varying functions, and an external forcing term. Our contributions can be summarized as follows. For the unforced problem, we established sharp conditions for finite-time blow-up and global existence, thereby extending the classical Fujita theory to the larger class of regularly varying functions. This provides a unified framework that recovers several earlier results as particular cases. For the forced problem, we proved nonexistence of global weak solutions under natural assumptions on the parameters and the external source $\mathbf{w}$. At the same time, we derived sufficient smallness conditions on both the initial data and the forcing term that guarantee the existence of global mild solutions. Our analysis combines semigroup estimates for the mixed local–nonlocal operator, test function techniques, and asymptotic properties of regularly varying functions, highlighting the interplay between diffusion mechanisms, temporal weights, and external forcing.
Despite these advances, several questions remain open and deserve further investigation. The long-time asymptotics of global solutions, such as decay rates, self-similar behavior, or convergence toward stationary states, remain largely unexplored in the present setting. Our approach could also be adapted to equations with gradient-type nonlinearities, coupled equation systems, or boundary value problems in bounded domains, where competition between local and nonlocal effects may lead to new phenomena. Finally, since the operator $\ell=\Delta-(-\Delta)^{\mathsf{s}}$ has deep connections with stochastic processes, it would be interesting to develop a probabilistic framework for our results, possibly linking blow-up behavior with properties of underlying Lévy-type processes.
Appendix A Regularly varying functions
For the sake of completeness, we present a brief overview of the principal properties of regularly varying functions.
The foundational results originate in Karamata’s seminal work [30] and de Haan’s thesis [18]. However, for the convenience of the reader, we refer primarily to the more accessible treatments available in the comprehensive monographs [10, 23, 19, 46], where these properties are systematically developed and rigorously presented.
Definition A.1.
A measurable function $\ell:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is said to be regularly varying at infinity with index $\rho\in\mathbb{R}$ if
$\lim_{\lambda\to\infty}\frac{\ell(\lambda\,x)}{\ell(\lambda)}=x^{\rho}\quad\text{for every }x>0.$ (A.1)
We denote this by $\ell\in\mathcal{RV}_{\rho}$. In the special case $\rho=0$, the function $\ell$ is called slowly varying at infinity (in the sense of Karamata). More precisely, $\ell$ is slowly varying at infinity if
$\frac{\ell(\lambda\,x)}{\ell(\lambda)}\to 1\quad\text{as }\lambda\to\infty,\quad\text{for all }x>0.$ (A.2)
Remark A.1.
  1. (i)
    The condition (A.2) captures the idea that $\ell$ varies very gradually at infinity.
  2. (ii)
    Slowly varying functions were first introduced by Karamata in [30, 31].
  3. (iii)
    If $\ell\in C^{1}$ near infinity, a sufficient condition for (A.2) to hold is
    $\lim_{x\to\infty}\frac{x\ell^{\prime}(x)}{\ell(x)}=0.$ (A.3)
One of the foundational results in the theory of regularly varying functions is the Uniform Convergence Theorem (UCT). First proved by Karamata in the continuous case and later extended to the measurable setting by Korevaar and collaborators in 1949. Given its importance, we state the theorem precisely below. For several proofs, see [10, Theorem 1.2.1, p. 6].
Theorem A.1.
If $\ell\in\mathcal{RV}_{\rho}$, then for arbitrarily chosen $a$ and $b$, where $0<a<b<\infty$, the equality (A.1) holds uniformly for $x\in[a,b]$.
Another fundamental result concerning slowly varying functions is their representation theorem, which plays a crucial role in various areas of analysis.
Theorem A.2 ([10, Theorem 1.3.1, p. 12]).
A measurable function $\ell$ is slowly varying if and only if it can be expressed in the form
$\ell(x)=c(x)\exp\left\{\int\limits_{a}^{x}\frac{\varepsilon(t)}{t}\,dt\right\}\quad(x\geq a),$ (A.4)
for some constant $a>0$, where $c(\cdot)$ is measurable with $c(x)\to c\in(0,\infty)$ and $\varepsilon(x)\to 0$ as $x\to\infty$.
Remark A.2.
  1. (i)
    Since $\ell$, $c$, and $\varepsilon$ may be modified freely on bounded intervals, the specific choice of the lower limit $a$ is not essential. For example, one may take $a=1$, or even $a=0$ by requiring $\varepsilon\equiv 0$ near the origin to ensure convergence of the integral. Moreover, the function $c$ can always be chosen to eventually be bounded.
  2. (ii)
    The representation (A.4) can be equivalently rewritten in the form
    $\ell(x)=\exp\left\{c_{1}(x)+\int\limits_{a}^{x}\frac{\varepsilon(t)}{t}\,dt\right\},$ (A.5)
    where $c_{1}(x)$ and $\varepsilon(x)$ are bounded measurable functions such that $c_{1}(x)\to d\in\mathbb{R}$ and $\varepsilon(x)\to 0$ as $x\to\infty$.
The representation formula (A.5) immediately yields the following asymptotic result. A proof can be found, for example, in [39].
Lemma A.1.
Let $\ell$ be a slowly varying function, $\alpha<0$ and $\beta\in\mathbb{R}$. Then
$x^{\alpha}\left(\ell(x)\right)^{\beta}\to 0\quad\text{as}\quad x\to\infty.$ (A.6)
Theorem A.3 (Karamata’s theorem for regularly varying functions [30, 18]).
Let $\ell:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a Lebesgue integrable function on every finite interval.
  1. (i)
    Suppose $\rho\geq-1$ and $\ell\in\mathcal{RV}_{\rho}$. Then
    $x\longmapsto\int\limits_{0}^{x}\ell(t)\,dt\in\mathcal{RV}_{\rho+1},$
    and
    $\lim_{x\to\infty}\left(\frac{x\ell(x)}{\int\limits_{0}^{x}\ell(t)\,dt}\right)=\rho+1.$ (A.7)
  2. (ii)
    Suppose $\rho<-1$ and $\ell\in\mathcal{RV}_{\rho}$. Then the tail integral
    $\int\limits_{x}^{\infty}\ell(t)\,dt<\infty,$
    and satisfies
    $\int\limits_{x}^{\infty}\ell(t)\,dt\in\mathcal{RV}_{\rho+1},$
    together with the asymptotic relation
    $\lim_{x\to\infty}\left(\frac{x\ell(x)}{\int\limits_{x}^{\infty}\ell(t)\,dt}\right)=-\rho-1.$ (A.8)
  3. (iii)
    Suppose $\rho=-1$ and $\int\limits_{x}^{\infty}\ell(t)\,dt<\infty$. Then the asymptotic identity (A.8) also holds.
Remark A.3.
We now offer an intuitive interpretation of the asymptotic equalities (A.7) and (A.8). Let $f\colon(0,\infty)\to(0,\infty)$ be a positive function that is Lebesgue integrable on every finite interval. Consider the following two cases:
  1. (i)
    Suppose $f(x)=\frac{\ell(x)}{x^{\alpha}}$, where $\alpha<1$, and $\ell$ is a slowly varying function at infinity. Then, as $x\to\infty$, we have
    $\int\limits_{0}^{x}f(t)\,dt=\int\limits_{0}^{x}\frac{\ell(t)}{t^{\alpha}}\,dt\sim\frac{\ell(x)}{(1-\alpha)x^{\alpha-1}}=\frac{xf(x)}{1-\alpha}.$ (A.9)
  2. (ii)
    Suppose instead that $f(x)=\frac{\ell(x)}{x^{\alpha}}$, where $\alpha>1$, and again $\ell$ is slowly varying at infinity. Then, as $x\to\infty$, we find
    $\int\limits_{x}^{\infty}f(t)\,dt=\int\limits_{x}^{\infty}\frac{\ell(t)}{t^{\alpha}}\,dt\sim\frac{\ell(x)}{(\alpha-1)x^{\alpha-1}}=\frac{xf(x)}{\alpha-1}.$ (A.10)
In both cases, the idea is that the asymptotic behavior of the integral can be captured by treating the slowly varying part $\ell(x)$ as approximately constant and integrating the dominant power-law component. This leads to a simple but useful approximation of the integral in terms of the original function $f(x)$.
In [14], the authors developed a generalized framework that extends the classical class $\mathcal{RV}_{\rho}$, allowing for functions whose asymptotic behavior resembles regular variation, even though the limit in (A.1) does not necessarily exist. More precisely, a first characterization of this new class is given below [14, Theorem 1.1, p. 111].
Definition A.2.
Consider a measurable function $U\colon(0,\infty)\to(0,\infty)$ that remains bounded on finite intervals. We say that $U$ belongs to the class $\mathcal{M}(\rho)$ if its logarithmic growth rate satisfies
$\lim_{x\to\infty}\frac{\log U(x)}{\log x}=\rho.$ (A.11)
Remark A.4.
  1. (i)
    If $U$ is slowly varying at infinity, then the limit in (A.11) holds with $\rho=0$. However, the converse does not hold in general. For instance, the function $U(x)=2+\sin x$ satisfies (A.11) with $\rho=0$, but it is not slowly varying due to its oscillatory behavior.
  2. (ii)
    It is shown in [14, Theorem 1.2, p. 111] that a function $U\in\mathcal{M}(\rho)$ if and only if it admits the representation
    $\ell(x)=\exp\left\{\alpha(x)+\int\limits_{a}^{x}\frac{\beta(t)}{t}\,dt\right\},\quad x\geq a>0,$ (A.12)
    where $\alpha(x)/\log x\to 0$ and $\beta(x)\to\rho$ as $x\to\infty$.
  3. (iii)
    The condition (A.11) captures functions whose asymptotic behavior mimics that of $x^{\rho}$, possibly modulated by a slowly varying function.
  4. (iv)
    Several illustrative examples of such functions are:
    $x^{\rho},\;\,\;x^{\rho}\,(\log x)^{\alpha},\;\,\;x^{\rho}\,\Biggl{(}1+\frac{\sin(\log x)}{\log x}\Biggr{)},\;\,\;x^{\rho}\,\Biggl{(}1+\frac{\sin(\log\log x)}{\log x}\Biggr{)}.$ (A.13)
  5. (v)
    Consider the function $U(x)=\exp(\sqrt{\log x}).$ Then,
    $\frac{\log U(x)}{\log x}=\frac{\sqrt{\log x}}{\log x}\to 0,$
    but this convergence is not of the form $\log x^{\rho}$, so $U\notin\mathcal{M}(\rho)$ for any $\rho\in\mathbb{R}$.
One of the characterization of $\mathcal{M}$ that will be useful for our purpose can be stated as follows.
Theorem A.4.
[14, Theorem 1.3] Let $U$ be a positive and measurable function with support $\mathbb{R}_{+}$ and bounded on finite intervals. Then $U\in\mathcal{M}(\rho)$ if and only if there exist slowly varying functions $\ell_{1}$ and $\ell_{2}$ such that
$\frac{U(x)}{x^{\rho}\ell_{1}(x)}\to 0\quad\text{and}\quad\frac{U(x)}{x^{\rho}\ell_{2}(x)}\to\infty\quad\text{as }x\to\infty.$ (A.14)
A key consequence of Theorem A.1, which will be utilized in deriving the Fujita exponent, is the following asymptotic estimate for integrals involving slowly varying functions.
Proposition A.1.
Let $\mathsf{h}:(0,\infty)\to(0,\infty)$ be a continuous function satisfying (1.10). Let $\beta\in\mathbb{R}$, and define
$F(R)=\int\limits\limits_{a}^{b}\left(\mathsf{h}(R\tau)\right)^{\beta}\,d\tau$ (A.15)
where $R>0$ and $0<a<b<\infty$. Then, as $R\to\infty$,
$F(R)\sim\left(\int\limits_{a}^{b}\tau^{\beta\gamma}\,d\tau\right)R^{\beta\gamma}\left(\ell(R)\right)^{\beta}.$ (A.16)
Proof.
From (1.10), we immediately obtain
$F(R)=R^{\beta\gamma}\int\limits_{a}^{b}\tau^{\beta\gamma}\,\ell(R\tau)^{\beta}\,d\tau.$
Since $\ell$ is slowly varying at infinity, Theorem A.1 implies that $\ell(R\tau)^{\beta}\sim\ell(R)^{\beta}$ uniformly for $\tau\in[a,b]$ as $R\to\infty$. Consequently,
$F(R)\sim R^{\beta\gamma}\left(\ell(R)\right)^{\beta}\int\limits_{a}^{b}\tau^{\beta\gamma}d\tau,$
which is the desired relation (A.16). The integral is finite since $0<a<b<\infty$. ∎
Declarations. On behalf of all authors, the corresponding author states that there is no conflict of interest. No data-sets were generated or analyzed during the current study.